OpenStudy (jade27p):

What is the average rate of change from x = –1 to x = 1?

OpenStudy (idkanything):

1 is average

OpenStudy (jade27p):

how

OpenStudy (jade27p):

i was thinking 2

Nnesha (nnesha):

the question is incomplete

OpenStudy (idkanything):

yes im sorry I was thinking adding what are the choices?

Nnesha (nnesha):

there should be a graph or a function average rate of change means slope first of all find y values when `x=-1 and x=1` then use the slope formula \[\huge\rm \frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (idkanything):

Good job Nnesha

OpenStudy (jade27p):

graph of parabola going through negative 1, 1, and 1, negative 3

OpenStudy (jade27p):

that's the graph

Nnesha (nnesha):

alright so order pair (-1,1) and (1,3) \[( \color{Red}{-1},\color{blue}{1})(\color{red}{1},\color{blue}{3}) \] \[\rm (\color{Red}{x_1},\color{blue}{y_1})(\color{reD}{x_2},\color{blue}{y_2})\]

Nnesha (nnesha):

substitute x's and y's with these values \[\large\rm \frac{ y_2- y_1 }{ x_2-x_1 }\] x_1=-1 x_2=1 y_1=1 y_2=3

OpenStudy (jade27p):

would the first be -1

OpenStudy (jade27p):

would the answer be 4

Nnesha (nnesha):

how did you get 4 ?

Nnesha (nnesha):

subtract y values subtract x values and then divide y/x what would you get ?

OpenStudy (jade27p):

i got 2 but it was wrong

OpenStudy (perl):

You should get -2. Did you follow Nnesha's work?

Nnesha (nnesha):

\[\huge\rm (\color{red}{-1},\color{blue}{1})(\color{orange}{1},\color{green}{-3})\]\[\huge\rm (\color{red}{x_1},\color{blue}{y_1})(\color{orange}{x_2},\color{green}{y_2}) \] \[\huge\rm \frac{\color{green}{ y_2}-\color{blue}{y_1} }{\color{orange}{ x_2}-\color{red}{x_1} }\] \[\huge\rm \frac{\color{green}{- 3}-\color{blue}{1} }{\color{orange}{ 1}-\color{red}{(-1)} }=\frac{-3-1}{1+1}=\]

Nnesha (nnesha):

ahh i see sorry i forgot the negative sign